OpenStudy (anonymous):

Solve 2x2 + 12x= 10 answers: -3+-√14 -3+-2√2 -3+-√19 -3+-√29

OpenStudy (konradzuse):

ok so lets start this out simply.

OpenStudy (anonymous):

ok

OpenStudy (konradzuse):

divide both sides by 2 so we have x^2 + 6x = 5

OpenStudy (anonymous):

okay.

OpenStudy (konradzuse):

now we are going to do something a little funky/interesting to get what we want. add 9 to both sides so we get x^2 + 6x + 9 = 14. With this we can get (x+3)^2 = 14 understand how I got that?

OpenStudy (anonymous):

yah.

OpenStudy (konradzuse):

now we will take the sqrt of each side to get |x+3| = sqrt(14)

OpenStudy (anonymous):

uh-huh

OpenStudy (konradzuse):

-3 from each side to get x = -3+- sqrt(14) so A

OpenStudy (anonymous):

ty :)

OpenStudy (konradzuse):

np :)